a) determine the area bounded by the graph of f(x)=3x2 and the x axis between x=0 and x=1

b) evaluate intergral \int \frac{2x-2}{x^{2}-2x}dx
Answer #1

a) The required area is given by

  3.2-2 di (by definition)

3x2 da = 2*6 = 1–0=1 5

b) 2.0 – 2 -do C2 - 2.0

Substitute ,2 T - 2

We have dt dr = 2.1 – 2 dt = 2.1 – 2 dr

Plug the values in the previous integral:

dt 2.0 – 2 - dac 22 2.c t

dt = In It + t

Back substitute to get

2.1 – 2 22 d.r = ln |22 – 2x + 200